Phenylcarbinol is used in nasal sprays as a preservative. A solution of 0.52 g of the compound in 25.0 g of water has a melting point of-0.36 °C. What is the molar mass of phenylcarbinol? (Kp for water is -1.858°C/m.)
1
Expert's answer
2019-05-22T01:00:07-0400
"n= \frac{m}{M}= \frac{0.52 g}{M}""molarity= \frac{n}{m(water)}=\frac{\frac{0.52 g}{M}}{0.025 kg}= \frac{0.52 g}{M\times0.025}"

"\Delta T_{fp}=k_f\times molarity"

"-0.36= -1.858\times \frac{0.52}{M\times 0.025}"

"M=107.35\frac{g}{mol}"